The Addison Company also purchased machinery on May 25, 2001 that had an original cost of $32,500 with an estimated salvage value of $2,500 and an estimated useful life of 10 years. What would be the book value using the straight-line method as of 12-31-04
Answer #1

Annual Depreciation Expense (Straignt LIne Method) = (Cost of Asset - Salvage Value) / Useful life

= (32,500 - 2,500) / 10

= $3,000

Year 2001 Depreciation Expense (221 days from 25 May to Dec 31) = $3,000 * 221 / 365 = $1,816

Year 2002 Depreciation Expenses = $3,000

Year 2003 Depreciation Expenses = $3,000

Year 2004 Depreciation Expense = $3,000

Total Accumulated Depreciation till 12-31-04 = $3,000*3 + $1,816 = $10,816

Book Value as of 12-31-04 = Cost of Asset - Accumulated Depreciation till 12-31-04

= $32,500 - $10,816

= $21,684

Hope the above calculations, working and explanations are clear to you and help you in understanding the concept of question.... please rate my answer...in case any doubt, post a comment and I will try to resolve the doubt ASAP…thank you